OpenStudy (sh3lsh):

A hydrogen atom in the ground state absorbs a photon of wavelength 95.0 nm. What energy level does the electron reach?

OpenStudy (sh3lsh):

I know the answer's 5. But can someone show me the calculations?

OpenStudy (anonymous):

\[E=\frac{ hc }{ \lambda }\]

OpenStudy (anonymous):

Ground energy Level of Hydrogen is -13.2eV

sam (.sam.):

Wavelength given is \(95.0 nm=95.0 \times 10^{-9}m\) \[\lambda=95 \times 10^{-9} \\ \\ E=\frac{hc}{\lambda} \\ \\ E=\frac{(6.63\times 10^{-34})(3 \times 10^8)}{(95\times 10^{-9})} \\ \\E=2.094 \times 10^{-18}J \\ \\ \text{For each energy level,} \\ \\ \frac{2.094\times 10^{-18}}{e}=13.1eV\] 13.1 Which is the fifth energy level.

OpenStudy (anonymous):

\[E=\frac{ 2.18*10^-18 Z^2 }{ n^2 }\]

OpenStudy (aaronq):

Use the rydberg equation: \[\frac{ 1 }{ \lambda} = R (\frac{1}{n _{1}^{2} }- \frac{ 1 }{ n _{2}^{2} })\] \[\frac{ 1 }{ 9.5x10^-8} = 1.097x10^7 (\frac{1}{1^{2} }- \frac{ 1 }{ n _{2}^{2} })\] isolate \[n _{2}\]

OpenStudy (aaronq):

ps. i messed up the order of the n's it should read: \[(\frac{ 1 }{ n _{f}^{2} }-\frac{ 1 }{ n _{i}^{2} })\] i= initial, f= final \[(\frac{ 1 }{ n _{f}^{2} }-\frac{1}{1^{2} })\]